A job was timed for 60 cycles and had an average of 1.9 minutes per piece. The performance rating was 90 percent, and workday allowances are 18 percent. Determine each of the following:


a. Observed time. (Round your answer to 1 decimal place.)


  Observed time minutes


b. Normal time. (Round your answer to 2 decimal places.)


  Normal time minutes


c. Standard time. (Round your answer to 2 decimal places.)


  Standard time minutes
Answer #1

A. Observed time(OT) = Average time = 1.9 minutes

B. Normal time = OT X performance rating = 1.9 * 0.9 = 1.71 minutes

C. Standard Time = ST = NT X AF

AF = Allowance factor = 1 / (1 - allowance) = 1 / (1 - 0.18) = 1.22

ST = 1.71 * 1.22 = 2.0862 minutes

Do give a positive rating if you think this answer helped